# Precise calculation of the EFT likelihood with primordial non-Gaussianities

### v1

@@ 4.4 The conditional likelihood
We are now exceedingly close to the final result, yet a few remaining preparatory steps must still to be completed. We denote the logarithms of the two likelihoods as $\wp[\delta]$ and $\wp[\delta_{g},\delta]$, and discard all field-independent terms. This yields
@@
  • We begin by examining these stochastic noise terms. It is evident that when $P_{\epsilon_{g}\epsilon_{m}}=0$, the saddle-point contributions to both the matter likelihood and the joint likelihood exactly coincide, implying that the conditional likelihood upon introducing PNG remains the same as in the previous one. In most forward modeling approaches, considering only the saddle-point contributions typically provides a reasonable approximation, as noted in [23]. However, for more accurate theoretical predictions, it becomes essential to incorporate higher-order noise contributions into the analysis [56]. These noise terms are of critical importance, as they are introduced to cancel the UV dependence of the loop diagrams [23]. Consequently, their inclusion in the computation bears a natural analogy to the construction of the renormalized action in quantum field theory (and these noise terms can be interpreted as counterterms) [38]. Similarly, we can generalize the calculation to encompass higher-order scenarios, analogous to the 1PI [40, 57] and 2PI [42, 58] effective action in quantum field theory. Furthermore, it is the cross stochasticity between the matter and galaxy fields that induces the inconsistency and lack of cancellation of their respective saddle-point contributions, resulting in nontrivial contributions to the conditional likelihood when PNG are incorporated. This indicates that even at the same perturbative order, we cannot directly cancel the identical perturbative terms between the two likelihoods during the computation. We concluded that considering higher-order noise terms introduces new contributions to the conditional likelihood.
@@
  • We next compare the contributions from the saddle points and around the saddle points. In our original expectation, the effective interaction term in the partition functions introduced by PNG is proportional to $\delta_{\rm in}^{3}$, as shown in eqs. (3.11) and (3.13), thereby yielding additional terms in the equations of motion eqs. (4.10), (4.11). Thus, this contribution will be reflected in the field dependence of the saddle-point solution, which in turn induces a corresponding dependence of the fields on the eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ and $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$, then resulting in a irreducible factor in both likelihoods. Interestingly, when computing the matter likelihood, although all the eigenvalues exhibit explicit field dependence, this dependence cancels out upon performing the Gaussian integration, leading to trivial additional contributions. We observe that even when ignoring the cross stochastic term $P_{\epsilon_{g}\epsilon_{m}}$, the contributions around the saddle points differ fundamentally between the two likelihoods: one exhibits field dependence, whereas the other does not. Thus, we argue that the precise calculation in the presence of PNG will also yield nontrivial results even only considering the lowest-order noise.
@@
  • Here, we describe how to incorporate more cosmological effects into our calculation. Clearly, these effects can be manifested in the corresponding terms within the partition functions. For instance, following the spirit of effective field theory, the implementation of PNG also necessitates adjustments to the galaxy bias expansion via the introduction of new effective expansion operators [13]. These adjustments will be reflected in the reformulation of the expression for $\delta_{g,\rm fwd}[\delta_{\rm in}]$. We emphasize that in our calculations, although a simplified treatment of $\delta_{g,\rm fwd}$ is applied, none of the computational steps depend on its specific form. Therefore, when we want to incorporate this effect, we can directly modify $\delta_{g,\rm fwd}[\delta_{\rm in}]$ via the desired bias expansion and proceed with the calculation. Alternatively, if we seek to incorporate higher-order PNG, we would need to expand the non-Gaussian primordial potential, as given in eq. (3.1), to higher orders. This generalization will only lead to higher-order vertices of $\delta_{\rm in}$ in the partition functions eqs. (3.11) and (3.13). Thus, we can still extract the additional terms associated with the saddle-point contributions by appropriately shifting the saddle-point solutions. Meanwhile, around the saddle points, these vertices will also result in modifications on $P_{\rm in}^{-1}$. Similarly, we can also incorporate higher-order galaxy stochastistics (correspond to higher-order terms of $J$ and $J_{g}$) and viscous effects (correspond to $JJ\delta_{\rm in}$ term [23]), among others. The detailed corrections to the conditional likelihood induced by these effects will be left for future work to investigate.
  • Our result may also provide new insights into Bayesian forward modeling [59, 60, 61, 62, 63] approaches. In fact, the Bayesian method is profoundly reliant on the conditional likelihood we obtained in this work, which makes the PNG corrections we have derived crucially important. At the same time, when we aim to incorporate any cosmological effects into this method, we only need to add the corresponding terms to the partition functions and then employ the corrected conditional likelihood. Conversely, the Bayesian framework can also serve to constrain the parameters that characterize these effects. Specifically, in this case, it facilities the constraining of parameters that govern the primordial bispectrum. Our results will contribute to the development of more comprehensive forward models, enhance the accuracy of parameter constraints, and increase the flexibility of the model.
@@ 5 Summary and conclusion
The conditional likelihood is the key physical quantity required when implementing Bayesian forward modeling. In this paper, we provide a functional formulation for the precise calculation of the EFT conditional likelihood, considering non-Gaussian initial conditions. The precision here is reflected in two aspects: first, we have considered the contribution of higher-order noise terms in the partition functions eqs. (3.11) and (3.13); second, we have not only accounted for the contributions from the saddle points but also incorporate contributions from their neighborhoods. And our results indicate that, although the conditional likelihood assuming only the lowest-order noise term via saddle-point approximation method is identical to that derived under Gaussian initial conditions (consistent with [23]), these two effects yield irreducible, field-dependent additional contributions. This suggests that, in the context of forward modeling, it is appropriate to introduce these additional terms corresponding to PNG within the conditional likelihood.
@@
In our calculation, we employed the saddle-point expansion method, which is important in the computation of decay rates. This method enables us to approximate the path integral by considering both the saddle-point contributions and the contributions from their neighborhoods. When evaluating the latter, it is necessary to consider the issue of the negative-mode solutions for the second derivative of the action at the saddle points. Should such solutions exist, we need to systematically deform the original integration contour into the sum of the steepest descent contours in the complexified field space, guaranteeing a real and convergent result. The application of the steepest descent method in the computation of the likelihoods is analogous to its use in quantum field theory with multiple fields [58], as clearly demonstrated in both [24] and the examples presented in this work. At the same time, we contend that the incorporation of higher-order noise terms is of significant importance, as their inclusion effectively cancels out the UV cutoff dependence inherent in the loop diagrams, analogous to the renormalization procedure in quantum field theory [38]. Therefore, we believe that the corrections to the non-Gaussian conditional likelihood presented in this paper are essential for deriving more stringent constraints on the parameters associated with PNG.
@@
Comparing the expressions for the likelihoods with and without the inclusion of PNG is also of great interest. After the implementation of leading-order PNG, identical additional terms are introduced into the partition functions of both matter and galaxy fields. However, once an additional source of noise is taken into account, the saddle-point contributions to the matter likelihood and joint likelihood integrals no longer cancel out, thereby justifying the involution of corresponding effective terms in simulations. Another noteworthy point is that, even in the absence of the cross stochasticities between the matter and galaxy fields, the results of the two likelihoods remain inconsistent around the saddle points. Therefore, we regard our results as both meaningful and of practical relevance.
@@
There remain unresolved issues in our work that cannot be addressed from a purely theoretical perspective. In our calculations, there are three parameters that are suppressed on large scales: the PNG parameter $f_{\rm NL}$, higher-order noise terms, and the higher-order expansions around the saddle points. We treat these parameters as perturbative quantities, whose higher-order contributions are progressively suppressed. However, we are currently unable to compare the relative magnitudes of these contributions, although certain parameters can be partially disentangled by performing a rescaling of the momentum. For instance, we may be interested in comparing the contributions around the saddle points with respect to the order of non-Gaussian perturbations, as this enables a more precise extraction of the relevant perturbative terms. These aspects require further investigation through simulations and observational data to determine, and Bayesian forward modeling will in return help us strengthen our theoretical framework.
@@ Acknowledgments
JYK thanks Minxing Li for his useful discussions. We acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).

### v3

@@ 4.4 The conditional likelihood
We are now exceedingly close to the final result, yet a few remaining preparatory steps must still be completed. We denote the logarithms of the two likelihoods as $\wp[\delta]$ and $\wp[\delta_{g},\delta]$, and discard all field-independent terms. This yields
@@
  • We begin by examining these stochastic noise terms. It is evident that when $P_{\epsilon_{g}\epsilon_{m}}=0$, the saddle-point contributions to both the matter likelihood and the joint likelihood exactly coincide, implying that the conditional likelihood upon introducing PNG remains the same as in the previous one. In most forward modeling approaches, considering only the saddle-point contributions typically provides a reasonable approximation, as noted in [23]. However, for more accurate theoretical predictions, it becomes essential to incorporate higher-order noise contributions into the analysis [26]. These noise terms are of critical importance, as they are introduced to cancel the UV dependence of the loop diagrams [23]. Consequently, their inclusion in the computation bears a natural analogy to the construction of the renormalized action in quantum field theory (and these noise terms can be interpreted as counterterms) [45]. Similarly, we can generalize the calculation to encompass higher-order scenarios, analogous to the 1PI [47, 60] and 2PI [49, 61] effective action in quantum field theory. Furthermore, it is the cross stochasticity between the matter and galaxy fields that induces the inconsistency and lack of cancellation of their respective saddle-point contributions, resulting in nontrivial contributions to the conditional likelihood when PNG are incorporated. This indicates that even at the same perturbative order, we cannot directly cancel the identical perturbative terms between the two likelihoods during the computation. We conclude that considering higher-order noise terms introduces new contributions to the conditional likelihood.
@@
  • We next compare the contributions from the saddle points and around the saddle points. In our original expectation, the effective interaction term in the partition functions introduced by PNG is proportional to $\delta_{\rm in}^{3}$, as shown in eqs. (3.11) and (3.13), thereby yielding additional terms in the equations of motion eqs. (4.10), (4.11). Thus, this contribution will be reflected in the field dependence of the saddle-point solution, which in turn induces a corresponding dependence of the fields on the eigenvalues of $S^{\prime\prime}[\bar{\boldsymbol{\phi}}]$ and $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$, resulting in an irreducible factor in both likelihoods. Interestingly, when computing the matter likelihood, although all the eigenvalues exhibit explicit field dependence, this dependence cancels out upon performing the Gaussian integration, leading to trivial additional contributions. We observe that even when ignoring the cross stochastic term $P_{\epsilon_{g}\epsilon_{m}}$, the contributions around the saddle points differ fundamentally between the two likelihoods: one exhibits field dependence, whereas the other does not. Thus, we argue that the precision calculation in the presence of PNG will also yield nontrivial results even when only considering the lowest-order noise.
@@
  • Here, we describe how to incorporate more cosmological effects into our calculation. Clearly, these effects can be manifested in the corresponding terms within the partition functions. For instance, following the spirit of effective field theory, the implementation of PNG also necessitates adjustments to the galaxy bias expansion via the introduction of new effective expansion operators [13]. These adjustments will be reflected in the reformulation of the expression for $\delta_{g,\rm fwd}[\delta_{\rm in}]$. We emphasize that in our calculations, although a simplified treatment of $\delta_{g,\rm fwd}$ is applied, none of the computational steps depend on its specific form. Therefore, when we want to incorporate this effect, we can directly modify $\delta_{g,\rm fwd}[\delta_{\rm in}]$ via the desired bias expansion and proceed with the calculation. Alternatively, if we seek to incorporate higher-order PNG, we would need to expand the non-Gaussian primordial potential, as given in eq. (3.1), to higher orders. This generalization will only lead to higher-order vertices of $\delta_{\rm in}$ in the partition functions eqs. (3.11) and (3.13). Thus, we can still extract the additional terms associated with the saddle-point contributions by appropriately shifting the saddle-point solutions. Meanwhile, around the saddle points, these vertices will also result in modifications to $P_{\rm in}^{-1}$. Similarly, we can also incorporate higher-order galaxy stochastics, which correspond to higher-order terms of $J$ and $J_{g}$, and viscous effects, which correspond to $JJ\delta_{\rm in}$ term [23], among others. The detailed corrections to the conditional likelihood induced by these effects will be left for future work to investigate.
4.5 Connections to Bayesian forward modeling
We conclude this section with a discussion of the implication of our results for Bayesian forward modeling [62, 63, 64, 65, 66]. In fact, Bayesian methods rely fundamentally on the conditional probability density functional $\mathcal{P}[\delta_{g}|\delta_{\rm in}]$ that describes the likelihood of observing the tracer field $\delta_{g}$, conditioned on a given realization of the initial conditions $\delta_{\rm in}$. According to [22, 23], we can generally decompose $\mathcal{P}[\delta_{g}|\delta_{\rm in}]$ into two quantities: i) the conditional likelihood $\mathcal{P}[\delta|\delta_{\rm in}]$, i.e. the probability of having a total (dark+luminous) matter field $\delta$ given the initial conditions $\delta_{\rm in}$. ii) the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$ of observing $\delta_{g}$ given a matter field $\delta$. The former is generally informed by forward models of the matter distribution (such as $N$-body simulations), wheres the latter often requires support form semi-analytical calculations. The aim of this work, along with that of several recent studies [22, 23, 24, 25], is to provide an analytical and precision computation of the latter.
In this work, we present the formal treatment of the impact of PNG on the conditional likelihood $\mathcal{P}[\delta_{g}|\delta]$, yielding two kinds of irreducible contributions in the final result: one arising from the saddle-point contributions, scaling as $k^{2}f_{\rm NL}$, and the other from fluctuations around the saddle points, which scale as $f_{\rm NL}$ and $k^{2}f_{\rm NL}$, respectively. However, it is important to note that a field-level analysis on PNG data necessarily requires incorporating non-Gaussian $\mathcal{P}[\delta|\delta_{\rm in}]$ in the first step, prior to applying the likelihood derived in this work [20, 21]. Consequently, the effects of PNG are already partially embedded at this stage. As a result, even if a likelihood without PNG is employed in the second step, the inferred value of $f_{\rm NL}$ may be biased, yet it will not trivially converge to zero. To estimate the potential impact of our results on parameter interference, we expand eq. (4.49) perturbatively with respect to $f_{\rm NL}$,
$\wp[\delta_{g}|\delta]$ $\approx-S[\bar{\boldsymbol{\phi}}]+S_{g}[\bar{\boldsymbol{\phi}}_{g}]-f_{\rm NL}B\left(\frac{\delta}{D_{1}}-\frac{iP_{\epsilon_{g}\epsilon_{m}}X_{g,\rm cl}}{D_{1}}\right)S_{g}[\bar{\boldsymbol{\phi}}_{g}]$
$\sim\text{Gaussian results}\,+\underset{\rm saddle-point\,fluctuations}{\mathcal{O}(f_{\rm NL})}+\underset{\rm saddle\,point+fluctuations}{\mathcal{O}(k^{2}f_{\rm NL})}+\cdots\,$, (4.50)
where $\cdots$’ denotes higher-order terms. Our analysis reveals that the leading contribution proportional to $f_{\rm NL}$, which arises from saddle-point fluctuations, dominates the corrections, and constitutes a unique prediction of the application of the saddle-point expansion method. On the other hand, the next-to-leading term, proportional to $k^{2}f_{\rm NL}$, is subject to a double suppression due to both its higher-derivative nature and the small amplitude of PNG. Nevertheless, recent PNG Bayesian forward modeling have begun to incorporate this order of contribution  [^8] This can be seen from the galaxy bias expansion present in [21]. to help us provide more precise constraints on PNG parameters. Accordingly, we choose to make the explicit expression of this order term manifest, in preparation for its potential relevance in future analyses. Compared to extracting primordial non-Gaussian signals from CMB anisotropies, probing such signatures directly from LSS is significantly more challenging due to non-linear fluctuations [67]. We hope our work contribute to the development of more comprehensive forward models, enhance the accuracy of parameter constraints, and increase the flexibility of the future simulations.
@@ 5 Summary and conclusion
The conditional likelihood is the key physical quantity required when implementing Bayesian forward modeling. In this paper, we provide a functional formulation for the precision calculation of the EFT conditional likelihood, considering non-Gaussian initial conditions. The precision here is reflected in two aspects: first, we have considered the contribution of higher-order noise terms in the partition functions eqs. (3.11) and (3.13); second, we have not only accounted for the contributions from the saddle points but also incorporated contributions from their neighborhoods. Our results indicate that, although the conditional likelihood assuming only the lowest-order noise term via saddle-point approximation method is identical to that derived under Gaussian initial conditions (consistent with [23]), these two effects yield irreducible, field-dependent additional contributions. This suggests that, in the context of forward modeling, it is appropriate to introduce these additional terms corresponding to PNG within the conditional likelihood.
@@
In our calculation, we employed the saddle-point expansion method, which is crucial in the computation of decay rates. This method enables us to approximate the path integral by considering both the saddle-point contributions and the contributions from their neighborhoods. When evaluating the latter, it is necessary to consider the issue of negative-mode solutions for the second derivative of the action at the saddle points. Should such solutions exist, we need to systematically deform the original integration contour into the sum of the steepest descent contours in the complexified field space, ensuring a real and convergent result. The application of the steepest descent method in the computation of the likelihoods is analogous to its use in quantum field theory with multiple fields [61], as clearly demonstrated in both [24] and the examples presented in this work. At the same time, we contend that the incorporation of higher-order noise terms is of significant importance, as their inclusion effectively cancels out the UV cutoff dependence inherent in the loop diagrams, analogous to the renormalization procedure in quantum field theory [45]. Therefore, we believe that the corrections to the non-Gaussian conditional likelihood presented in this paper are essential for deriving more stringent constraints on the parameters associated with PNG.
@@
Comparing the expressions for the likelihoods with and without the inclusion of PNG is also of great interest. After the implementation of leading-order PNG, identical additional terms are introduced into the partition functions of both the matter and galaxy fields. However, once an additional source of noise is taken into account, the saddle-point contributions to the matter likelihood and joint likelihood integrals no longer cancel out, thereby justifying the inclusion of corresponding effective terms in simulations. Another noteworthy point is that, even in the absence of cross stochasticities between the matter and galaxy fields, the results of the two likelihoods remain inconsistent around the saddle points. Therefore, we regard our results as both meaningful and of practical relevance.
@@
There remain unresolved issues in our work that cannot be addressed from a purely theoretical perspective. In our calculations, three parameters are suppressed on large scales: the PNG parameter $f_{\rm NL}$, higher-order noise terms, and higher-order expansions around the saddle points. We treat these parameters as perturbative quantities, whose higher-order contributions are progressively suppressed. However, we are currently unable to compare the relative magnitudes of these contributions, although certain parameters can be partially disentangled by performing a rescaling of the momentum. For instance, we may be interested in comparing the contributions around the saddle points with respect to the order of non-Gaussian perturbations, as this enables a more precise extraction of the relevant perturbative terms. These aspects require further investigation through simulations and observational data, and Bayesian forward modeling will, in return, help us strengthen our theoretical framework.
@@ Acknowledgments
JYK thanks Minxing Li for his useful discussions. We thank the anonymous reviewer for his/her professional and invaluable comments and suggestions. We acknowledge the support by the National Science Foundation of China (No. 12147217, 12347163), the China Postdoctoral Science Foundation (No. 2024M761110), and the Natural Science Foundation of Jilin Province, China (No. 20180101228JC).